OpenStudy (diamond123):

can anyone tell me how to use prime factorization when reducing multiplied fractions

OpenStudy (anonymous):

18/24 18 = 2*3*3 24= 2*2*2*3 There is a 2 and a 3 common to both so u can cancel with 6 3/4